### v1

# Busting_the_princelings_replication.bib Are Princelings Truly Busted? Evaluating Transaction Discounts in China’s Land Market

Abstract
This paper narrowly replicates Chen and Kung’s 2019 paper ( 134(1): 185-226). Inspecting the data reveals that nearly one-third of the transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number. Replicating the analysis on the data sans-duplicates yields a slightly smaller but still statistically significant princeling effect, robust across the regression results. Further analysis also reveals that coefficients interpreted as the effect of logarithm of area actually reflect the effect of scaled values of area; this paper also reinterprets and contextualizes these results in light of the true scaled values.
Keywords— data reproducibility, empirical validation, institutional reform, Chinese real estate
1 Introduction
The recent implosion of several leading Chinese real estate developers, such as Evergrande and Country Garden, has kindled interest in the mechanisms of the Chinese real estate market. Unique not only for its highly leveraged financing practices, the market is also a nexus of interactions between state-owned and private enterprises and has historically been plagued by corruption and cronyism, particularly on the land sale side. One often-cited paper on the topic is Chen and Kung (2019), titled “Busting the ‘Princelings’: The Campaign against Corruption in China’s Primary Land Market,” which explores the unique mechanism of Chinese cronyism in the land sale market. Specifically, they examine how “princeling” firms—those firms tied to family members of China’s top governing body, the Politburo—receive substantial price discounts on land parcels after controlling for a variety of location-based and transaction-level variables ([chen_busting_2019]). They also have significant findings relating princeling power, promotion likelihood, and magnitude of discount: more powerful princelings obtain larger discounts, and provincial party secretaries providing discounts were more likely to be promoted, with the likelihood increasing by the magnitude of the discount and the area of discounted land sold.
Between containing detailed information about officials’ promotions and establishing firms’ state-ownership and princeling statuses, the paper’s replication data is also of popular interest and has over 11,600 downloads from its archive on the Harvard Dataverse.111The paper’s replication data can be found at https://doi.org/10.7910/DVN/XW6OJT.
This paper replicates the results of Chen and Kung (2019, hereafter “CK”), accounting for cases of apparent duplicates and verifying data consistency. An initial inspection of CK’s data reveals that nearly one-third of the transactions (388,903 out of 1,208,621) are perfect duplicates of other rows, excluding the transaction number: these duplicates suggest, for instance, that the same buyer purchased two (or more) identical parcels of land with the same quality rating, area, and usage in the same city, month, and year, at the same price. Returning to the initial database and examining records of land sale transactions suggests that some of these duplicates correctly refer to identical yet distinct parcels sold at the same time, but many other duplicates erroneously reference the same parcel of land multiple times. Importantly, over one-half of the princeling parcels in the full dataset (9,815 out of 19,812) are duplicates. Replicating the analysis on the data, sans-duplicates, yields similar results to CK’s original findings: while the magnitude of discount for princelings is slightly smaller than that found using the original dataset, the princeling effect is still large and statistically significant across the regression results.
Yet, further inspection reveals that when CK aggregate parcel areas to group parcels (for instance, summing data from transactions to calculate the total area of land sold to each firm in each year), these area measures are misinterpreted: while CK interpret the variable as the logarithm of area, it is actually area divided by 1,000,000. I thus reinterpret CK’s originally calculated coefficients as scaled area values rather than logged area values, with and without duplicates. While the princeling effect remains stable, its magnitude becomes quite large, suggesting that princelings on average purchase tens of thousands of additional square meters of land per year compared to their nonprinceling peers—areas double, triple, or even quadruple the median parcel size, depending on the specification. In the Appendix, I include calculations utilizing the corrected logarithm of area purchased; these results are strange and somewhat unstable, with the princeling effect not only changing sign and direction between tables, but also dropping sharply prior to the anticorruption campaign.
2 Apparent Duplicates in the Data
In their paper, CK construct a set of “princeling” firms from several online sources and match them to land transaction data obtained from the Ministry of Land and Resources. I focus on this latter data, extracted from the Land Transaction Monitoring System (www.landchina.com). As each municipality’s “bureau of land and resources is required to report each land transaction in their jurisdiction electronically on this website,” per the Law of Land Management (p. 199), the dataset captures all land transactions.222China eliminated the secondary market for land use rights (LURs) in 2004, meaning the State is the sole seller of land, doing so at the municipal level ([li_urban_2019]; [gyourko_land_2022]). The Land Transaction Monitoring System provides details on land transactions with increasing specificity in recent years, detailing everything from transaction details like total payment amount, sale method, buyer name, and total area of land sold, to building restrictions, like floor area ratio and greening ratio, to location features of the parcel, including its province, city, and district, as well as a qualitative description of its location.333CK note that the system also provides the address and postal code of the parcel, “a two-digit code of land usage… [and] a three-digit industry code of the buyer’s firm” (p. 200). These fields are no longer included on the Land Transaction Monitoring System website nor are they included in CK’s replication data, which makes determining the extent of duplicates more complicated; it is possible that these information fields were removed from the database sometime after CK collected the data. Alternatively, CK may have obtained a slightly more comprehensive version of the website’s data directly from the ministry.
Examining CK’s dataset reveals several cases of apparent duplicates: specifically, after removing the transaction identifier, which was likely added by CK in data processing, 388,903 land transactions appear to be perfect duplicates. This suggests, in effect, that in the same month and year, a single buyer purchased identical parcels for the same price and use in the same city. Table LABEL:table_of_duplicates shows the extent and distribution of these duplicates. Of the 1,208,621 land transactions in CK’s data, 819,718 are unique, and 299,167 of these have at least one duplicate. While most transactions have only one duplicate, others have several—as many as 158.
Importantly, a considerable number of princeling firms are represented among the duplicates: of the 19,812 transactions that were conducted by princelings in the original data, over half of these are duplicates, with a significant share having three or more seemingly identical transactions.
| Observations | |||
| Frequency of occurrence | All | State-owned | Princeling |
| 1 | 520,551 | 181,571 | 4,414 |
| 2 | 218,001 | 75,640 | 3,479 |
| 3 | 79,366 | 27,473 | 1,554 |
| 4 | 581 | 193 | 81 |
| 5-10 | 966 | 319 | 201 |
| 10-20 | 167 | 59 | 48 |
| 20-30 | 48 | 10 | 21 |
| 40-158 | 38 | 15 | 17 |
| Number of unique rows | 819,718 | 285,280 | 9,815 |
| Number of total rows in Chen and Kung | 1,208,621 | 420,004 | 19,812 |
| Number of duplicate rows | 388,903 | 134,724 | 9,997 |
| Number of unique rows with duplicates | 299,167 | 103,709 | 5,401 |
In light of these results, there are a couple of possibilities: first, that there are truly identical neighboring parcels that are sold at the same time and price to a single buyer. In effect, the Ministry of Land could be subdividing a larger parcel and selling it as two (or more) smaller parcels. Specific usages could also dictate purchasing multiple nearby parcels of the same size and usage at the same time—for instance, land for building mobile phone towers.444Investigating the Ministry of Land’s database illustrates that this is in fact the case: telecommunications companies like China Mobile have purchased several parcels of the same area at the same price in the same city—in this case, the parcels were quite small (48 square meters), likely for telecommunications equipment. Second, it is possible that the duplicate rows are erroneous and that the database itself has multiple land transactions that refer to the same underlying land parcel.
Examining CK’s replication data is not enough to confirm whether these duplicates refer to the same underlying land parcel, given they do not include each parcel’s specific address. Yet, the transaction details can be matched with the original data from the Ministry of Land’s database, and the additional information available on the website can be used to determine if multiple transaction listings refer to the same parcel. Specifically, after being identified in the database, each land transaction can be linked with a parcel number via the transaction’s original sale announcement (the “Transfer Announcement”), and one can confirm how many parcels of each size are up for sale in the area at the time. With this information, duplicate land transaction listings that refer to the same underlying parcel can be confirmed as erroneous duplicates.
Looking at the detailed transaction records indicates that both hypotheses are true: while there are some cases where identical parcels were sold to a single buyer with the same features, many other cases appear to refer to one underlying parcel and are incorrect duplicates. Some of these are quite easy to recognize as duplicate listings of the same land transaction when referencing the Ministry of Land’s database. For instance, inspecting the database shows multiple cases where two separate transaction listings have the same electronic lookup number—a strong indicator of an incorrect duplicate.555The electronic lookup number is the database’s main identifier/search key for unique land transactions. Having two “separate” transactions with the same electronic lookup number is thus impossible; it can only be one transaction listed twice in error. Examining the Transfer Announcement for these parcels further confirms that only one parcel of the given size is listed for sale. Thus, there should only be one listing for the parcel in the Ministry of Land’s database—and in CK’s data.
Other cases are slightly less clear but are still strongly indicative of incorrect duplicates. For instance, two transactions have different electronic lookup numbers, but investigating the Transfer Announcement suggests that both listings refer to the same underlying parcel.666 In many cases, examining the full transaction details of the two listings reveals minor discrepancies between them—not contradictions, but rather missing information. These omissions are non-essential details (e.g., greening ratio, floor area ratio) rather than key identifiers (e.g., buyer, area, quality). A second listing of the same parcel may thus exist simply to complete the initial record. The cases with extremely high numbers of duplicates appear to largely be erroneous; for instance, there are 73 duplicate entries of a parcel with an area of only 100 square meters purchased by a single buyer at the same price in the same area with the same usage, all marked as princelings. Another case with 71 apparent duplicates has a larger area, but in the online database, there is no record of even one transaction occurring with the given parcel features.
Additional cases appear to be partially correct, partially incorrect duplicates. For instance, the Ministry of Land’s database reveals two identical parcels were sold at the same time to a single buyer at the same price. Yet, CK’s dataset has three transactions that match these features rather than two.777This may reflect an erroneous duplicate entry being removed from the website after CK obtained their data. A sample of several such cases is included in the Appendix.
Further, some transactions appear identical except for the location dummy variables (near500 and near1500) created by CK. These are dummy variables that reflect whether a given parcel is within 500 or 1500 meters, respectively, of a parcel purchased by a princeling firm. Presumably, CK used the specific location data (not included in their replication files) to determine these binary values, and thus, having different binaries for these values implies that the parcels are indeed distinct. These rows are therefore not classified as duplicates in this paper.888There are approximately 100,000 of these rows in the data. I was unable to verify whether these parcels are distinct using the Ministry of Land’s database alone; the specific address information would be needed to ensure these transactions reflect separate parcels. Even if one wanted to remove these rows as duplicates, it is not apparent which observation contains the correct near500 and near1500 values without the parcel addresses. I thus leave these quasi-duplicates in, but note that their status as unique parcels should be verified further.
Thus, a matched comparison with the Ministry of Land’s database, tracing parcel features and Transfer Announcements, suggests that a significant portion of the duplicate rows are erroneous, referring to the same underlying transaction. Yet, in the absence of specific location details for the parcels, it is not possible to examine the exact extent of incorrect duplicates versus those that refer to different but identical parcels.
2.1 Replicating results without duplicates: Tables V and X
When replicating CK’s regression results, I drop all duplicate rows but recognize that this path excludes duplicate entries that refer to different but identical parcels. However, as highlighted in Table 2, the regression results are largely still similar for their main specification (CK’s Table V, “Princeling Purchase and Land Price, 2004-2016”),
As detailed by CK, the dependent variable “is the log of price (yuan per square meter) for land parcel sold by local government to firm for usage in month-year . The key explanatory variable, , is a dummy variable equal to 1 if a firm is connected to a princeling” and 0 otherwise (p. 203). They control for several transaction-level variables including land quality (as evaluated by the municipal government, on a scale of 1-20), state-ownership status, sale method, firm size, and the log(area/100), where area is the land area sold in square meters. They also have city-year-usage, industry, and month fixed effects, with standard errors clustered at the province and firm levels and singletons included.
Column (1) illustrates regression (1) run on the full dataset, while columns (2) and (3) are conducted on the matched sample of parcels within a specified radius (1,500 or 500 meters, respectively) from a princeling parcel. Columns (4), (5), and (6) repeat these regressions but now include princeling firms interacted with the dummy variable (Politburo Standing Committee Member), aiming to explore whether firms connected to these committee members received larger discounts. Likewise, columns (7), (8), and (9) include princeling firms interacted with a dummy variable that reflects whether a transaction occurred after the connected official was retired. In all of these cases, the princeling effect is strongly negative with small standard errors—as was the case in CK’s original results—suggesting that the princeling effect is robust
| Log of land price | |||||||||
| All | meters | meters | All | meters | meters | All | meters | meters | |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | |
| Original data | |||||||||
| Princeling firms | -0.808*** | -0.904*** | -0.844*** | -0.545*** | -0.666*** | -0.620*** | -0.808*** | -0.894*** | -0.835*** |
| (0.025) | (0.034) | (0.033) | (0.035) | (0.043) | (0.043) | (0.030) | (0.040) | (0.038) | |
| Princeling firms*PSCM | -0.442*** | -0.420*** | -0.396*** | ||||||
| (0.037) | (0.048) | (0.049) | |||||||
| Princeling firms*Retired | -0.001 | -0.051 | -0.044 | ||||||
| (0.056) | (0.063) | (0.058) | |||||||
| Number of observations | 1,144,507 | 334,232 | 191,896 | 1,144,507 | 334,232 | 191,896 | 1,144,507 | 334,232 | 191,896 |
| Adjusted | 0.692 | 0.727 | 0.755 | 0.692 | 0.728 | 0.756 | 0.692 | 0.727 | 0.755 |
| Data without duplicates | |||||||||
| Princeling firms | -0.750*** | -0.849*** | -0.804*** | -0.485*** | -0.602*** | -0.574*** | -0.737*** | -0.830*** | -0.789*** |
| (0.027) | (0.033) | (0.033) | (0.035) | (0.039) | (0.041) | (0.033) | (0.038) | (0.037) | |
| Princeling firms*PSCM | -0.488*** | -0.470*** | -0.436*** | ||||||
| (0.040) | (0.048) | (0.048) | |||||||
| Princeling firms*Retired | -0.063 | -0.090 | -0.069 | ||||||
| (0.057) | (0.062) | (0.056) | |||||||
| Number of observations | 779,372 | 248,627 | 136,480 | 779,372 | 248,627 | 136,480 | 779,372 | 248,627 | 136,480 |
| Adjusted | 0.693 | 0.716 | 0.736 | 0.694 | 0.717 | 0.737 | 0.693 | 0.716 | 0.736 |
despite the apparent duplicates. Interestingly, when duplicates are removed, the princeling effect itself appears slightly smaller; for instance, the price advantage for princeling-connected firms, based on results for the 500-meter-radius sample, is now a 55.3% discount compared to CK’s original 57% discount. However, the interacted terms (princeling firms with Politburo Standing Committee connections and princeling firms connected to retired officials) are consistently larger than the original data’s results.999Even when replicating these main regression specifications on a subset without duplicates—and then going further to exclude rows which have different values in the near500 and near1500 binaries but are otherwise duplicates—the regression results shrink slightly more towards zero but remain largely the same. Standard errors likewise remain similar.
CK also use this granular transaction data in their Table X, “Princeling Purchases and Land Prices after Xi Took Office,” which examines whether President Xi’s anticorruption campaign impacted the price discount given to princeling firms. I regenerate this table without duplicate rows (Table 3), and the results are again quite similar to CK’s original findings: for the 500-meter matched sample, CK found that princeling firms had previously obtained an average price discount of 60.1% that dropped by 11.7% after the introduction of President Xi’s anticorruption
| Log of land price | ||||||||||
| All | meters | All | meters | All | meters | All | meters | All | meters | |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | |
| Original data | ||||||||||
| Princeling firms | -0.907*** | -0.920*** | -0.825*** | -0.858*** | -0.870*** | -0.896*** | -0.907*** | -0.920*** | -0.818*** | -0.847*** |
| (0.029) | (0.040) | (0.024) | (0.032) | (0.028) | (0.035) | (0.029) | (0.040) | (0.023) | (0.028) | |
| Princeling firms | 0.318*** | 0.257*** | 0.140* | 0.093 | ||||||
| Transaction after 2012 | (0.047) | (0.058) | (0.052) | (0.054) | ||||||
| Princeling firms | 0.819*** | 0.695*** | 0.504*** | 0.420*** | ||||||
| Central inspection | (0.124) | (0.139) | (0.079) | (0.096) | ||||||
| Princeling firms | 0.614*** | 0.572*** | 0.449*** | 0.447*** | ||||||
| Xi-appointed officials | (0.055) | (0.051) | (0.064) | (0.059) | ||||||
| Princeling firms | 0.109 | 0.037 | ||||||||
| Pre-2012 inspection | (0.074) | (0.070) | ||||||||
| Number of observations | 1,144,507 | 191,896 | 1,144,507 | 191,896 | 1,144,507 | 191,896 | 1,144,507 | 191,896 | 1,144,507 | 191,896 |
| Adjusted | 0.692 | 0.755 | 0.692 | 0.755 | 0.692 | 0.755 | 0.692 | 0.756 | 0.692 | 0.755 |
| Data without duplicates | ||||||||||
| Princeling firms | -0.899*** | -0.923*** | -0.774*** | -0.823*** | -0.844*** | -0.884*** | -0.900*** | -0.923*** | -0.752*** | -0.802*** |
| (0.033) | (0.041) | (0.025) | (0.031) | (0.034) | (0.037) | (0.033) | (0.041) | (0.025) | (0.028) | |
| Princeling firms | 0.351*** | 0.298*** | 0.151** | 0.113* | ||||||
| Transaction after 2012 | (0.047) | (0.059) | (0.051) | (0.053) | ||||||
| Princeling firms | 0.766*** | 0.675*** | 0.488*** | 0.415*** | ||||||
| Central inspection | (0.115) | (0.135) | (0.081) | (0.099) | ||||||
| Princeling firms | 0.595*** | 0.571*** | 0.441*** | 0.445*** | ||||||
| Xi-appointed officials | (0.055) | (0.054) | (0.065) | (0.060) | ||||||
| Princeling firms | 0.028 | -0.025 | ||||||||
| Pre-2012 inspection | (0.077) | (0.073) | ||||||||
| Number of observations | 779,372 | 136,480 | 779,372 | 136,480 | 779,372 | 136,480 | 779,372 | 136,480 | 779,372 | 136,480 |
| Adjusted | 0.694 | 0.736 | 0.693 | 0.736 | 0.694 | 0.737 | 0.694 | 0.737 | 0.693 | 0.736 |
campaign (column (2)); without duplicates, the 60.2% discount before the anticorruption campaign drops by 13.8% after the campaign begins (Table 3, column (2)). The other columns are equally similar and again suggest the anticorruption campaign strongly impacts the magnitude of discounts obtained by princelings; these remaining columns explore changes in princeling advantage in the presence of interacted dummy variables involving different dimensions of the anticorruption campaign (e.g., whether a central inspection occurred in the province at the time of sale, whether the transaction occurred after the anticorruption campaign began, etc.). As with CK’s Table V, these results are robust even in the presence of potentially erroneous duplicate rows.
3 Misspecifying Area
For their later tables, CK filter and transform the land transaction records to group by relevant features. When they aggregate transaction records at the firm and year level—such that they list the total area of all of firm ’s transactions in a given year—the value that they define as the natural logarithm of the area purchased is actually the area purchased divided by 1,000,000. However, CK interpret these area/1,000,000 values as logarithms; I include specific examples in the Appendix. Thus, I re-interpret their coefficients for the tables impacted by this transformation error, for cases with and without duplicates. This error extends to CK’s Tables VI, VIII, IX, XI, and XII as well as Figure VI.101010The tables discussed previously (CK’s Tables V and X) do not suffer from these issues because they do not contain the aggregated measures of the logarithm of area, remaining at the parcel-level. The remaining tables (I, II, III, IV, and VII) offer summary information and are not impacted by the misinterpretation; further, of CK’s figures, only Figure VI is impacted by the interpretation error.
3.1 Replicating CK’s Table VI
CK’s Table VI “Quantity of Land Purchased by Princeling Firms, 2004-2016” regresses the area of land purchased in the primary land market, divided by 1,000,000, on the dummy indicator signaling a princeling connection, replicated in Panel A of Table 4.111111While CK state that industry fixed effects are included in Table VI, they are omitted from the calculation. While CK originally find that princeling-connected firms purchased 0.2% more land each year than their nonprinceling counterparts, the regression coefficient actually suggests that princeling-connected firms purchased an additional 2,000 square meters of land annually compared to their nonprinceling peers, across all land transactions combined. Contextually, 90.8% of the rows have a total annual area purchased of zero square meters, as few firms (only 14%) purchase land in multiple years between 2004 and 2016. Excluding the rows in the panel that have zero area, the median area of land purchased annually by a firm is 26,668 square meters; the mean is 68,193 square meters, as a small number of firms have very high yearly transaction amounts (up to 14.67 million square meters). I thus find that this median more accurately reflects the average firm’s yearly transaction volume. When the rows with zero area are included, the median is 0 square meters, and the mean is 6,250 square meters. Excluding zeros, the average number of parcels purchased annually per firm is 1, although the maximum is 826 parcel purchases.
The 2,000 square meter advantage is therefore a 7.5% increase over the median parcel size (excluding rows with zero area), much more than CK’s original 0.2%. In column (2), firms connected to Politburo Standing Committee members purchase 32,000 square meters of additional land per year compared to nonprincelings—more than doubling the median parcel size; this is far above the 3% that CK originally obtained (column (2), exp(0.001+0.031)). Column (3) indicates that retirement has a negligible effect on the quantity of land purchased.
When Table VI is recalculated without duplicates, as highlighted in Panel B of Table 4, the result of column (1) does not change, but column (2) now suggests that princelings with Politburo Standing Committee connections purchase 41,000 additional square meters of land a year than their nonprinceling peers. The retirement of an official (column (3)) is now associated with a significant increase in the area of land purchased by princelings, with princeling firms connected to retired officials purchasing, on average, an additional 59,000 square meters of land—bringing their total purchases to more than three times the median parcel size.121212Note that a relatively small number of firms are connected to retired officials (668), and the data suggests that among purchases made by firms with connections to retired officials, smaller parcels are disproportionately represented among the duplicates, causing this large increase in the coefficient. Of the unique transactions involving retired officials, 46% have at least one duplicate, with some having as many as 22 duplicates. Indeed, the mean and median yearly parcel purchase areas are significantly higher for firms connected to retired officials while those for nonprincelings and for princelings without this connection are roughly equal to each other.
| Area of land purchased/1,000,000 | ||||||
| Panel A: Original data | Panel B: Data without duplicates | |||||
| (1) | (2) | (3) | (1) | (2) | (3) | |
| Princeling firms | 0.002*** | 0.001** | 0.002*** | 0.002*** | -0.000 | 0.001** |
| (0.000) | (0.000) | (0.000) | (0.000) | (0.000) | (0.000) | |
| Princeling firms*PSCM | 0.031*** | 0.041*** | ||||
| (0.004) | (0.004) | |||||
| Princeling firms*Retired | -0.001 | 0.058*** | ||||
| (0.001) | (0.004) | |||||
| Adjusted | 0.015 | 0.016 | 0.015 | 0.009 | 0.010 | 0.010 |
3.2 Tables VIII, IX, and XII
For Tables VIII, IX, and XII, I am unable to replicate CK’s aggregated area measures. These specifications focus on the effect of business favor on promotion, examining whether there is an association between officials selling land to princelings for discounted prices and their likelihood of promotion. As such, they regress political turnover for provincial officials on different measures of princeling connections (Table VIII), political turnover for municipal officials (Table IX), and then political turnover for both types of officials (Table XII); they thus aggregate the total land sold to firms under each official’s purview by year and province/municipality, respectively. CK likely have another variable that connects a transaction to an official at the province/municipality level, but they exclude this linking variable from the replication data, preventing recalculation. Yet, in all cases, the logarithm of the area of land sold seems extremely small, suggesting that the “logarithm of area sold” is again area/1,000,000.131313For Table XII, CK utilize the logarithm of area as the independent variable for a set of regressions, calling it , yet based on the description of this variable as measuring the “quantity of land purchased,” it is not clear whether they mean to interpret it as the logarithm of area or as a scaled measure of area (p. 220). Regardless, as CK do not offer a detailed interpretation of these coefficients, the substantive implications do not change: the effect of on the promotion of party secretaries drops for transactions occurring after 2012 for provincial and municipal secretaries and in regions with central inspections for provincial party secretaries.
3.3 Table XI and Figure VI
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign.141414Note that one of Xi’s first initiatives when taking over as party leader was the anticorruption campaign—he announced at the 18th National Congress (the top meeting for setting agendas within the party) that he would work to catch the “tigers” (high-ranking corrupt officials) and “flies” (low-ranking corrupt officials) to ensure the party’s survival ([wong_new_2012]; [xi_power_2013]; [yuen_disciplining_2014]). Table XI thus regresses the land area purchased/1,000,000 on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurs in a province wherein Xi replaced the party secretary; these results are illustrated in Table 5. CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1%, but this advantage dropped by 2.3% after 2012. Reinterpreting these coefficients yields that princeling firms annually purchased 78,000 more square meters of land than their nonprinceling counterparts, on average, before 2012, dropping to 56,000 square meters afterwards. With the median total purchased by a firm in a given year and province being 26,114 square meters, these purchase advantages reflect a sizeable princeling benefit, with the princeling advantage suggesting that princelings purchased almost quadruple the median land purchase volume before 2012 and then only triple the median after 2012.151515The panel data used in Table XI lists each firm’s parcel purchases grouped by each province and year between 1991-2016. Given few firms make frequent land purchases, 95.4% of the observations have zero square meters of land purchased in a given province/year. The reported median includes only those nonzero transactions as the median would be 0 otherwise. The mean is 3,076 square meters when including rows with zero transactions per year/province and 66,423 square meters when these zeros are excluded.
| Area of land purchased/1,000,000 | ||||||||
| Panel A: Original data | Panel B: Data without duplicates | |||||||
| (1) | (2) | (3) | (4) | (1) | (2) | (3) | (4) | |
| Princeling firms | 0.078*** | 0.073*** | 0.075*** | 0.078*** | 0.048*** | 0.050*** | 0.050*** | 0.048*** |
| (0.008) | (0.007) | (0.007) | (0.008) | (0.005) | (0.005) | (0.005) | (0.005) | |
| Princeling firms | -0.022** | -0.015 | 0.006 | 0.013 | ||||
| Transactions after 2012 | (0.006) | (0.007) | (0.005) | (0.007) | ||||
| Central inspection | 0.038*** | 0.025*** | 0.038*** | 0.025*** | ||||
| (0.002) | (0.004) | (0.001) | (0.004) | |||||
| Princeling firms | -0.053*** | -0.025*** | -0.032*** | -0.024** | ||||
| Central inspection | (0.007) | (0.007) | (0.006) | (0.007) | ||||
| Xi-appointed officials | 0.036*** | 0.028*** | 0.036*** | 0.029*** | ||||
| (0.002) | (0.002) | (0.002) | (0.003) | |||||
| Princeling firms | -0.056*** | -0.036*** | -0.033*** | -0.035*** | ||||
| Xi-appointed officials | (0.006) | (0.006) | (0.005) | (0.006) | ||||
| Adjusted | 0.032 | 0.047 | 0.052 | 0.057 | 0.021 | 0.038 | 0.044 | 0.050 |
In column (4), CK originally find that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant when the dummy variables for central inspection and Xi-appointed officials are included. They further note that “given that the sum of the coefficients in column (4) is not significantly different from zero, these two measures together effectively eliminate the advantage of the princeling firms in purchasing a larger quantity of land” (p. 219).161616Here, CK are summing the coefficients that include —that is, the dummy variable and its interaction terms. This result remains in the absence of the intended logarithm.
When duplicates are removed (Table 5, Panel B), the magnitude of the princeling effect drops across all specifications but remains statistically significant: column (1), for instance, indicates that princeling firms purchased 48,000 square meters more land than their nonprinceling counterparts annually. After 2012, this effect does not change in a statistically significant way. For column (4), as above, “transactions after 2012” is insignificant when “central inspection” and “Xi-appointed officials” are included, the coefficient changing sign now that duplicates have been removed, and the sum of the princeling dummy variable and its interacted terms is not statistically different from zero. “Transactions after 2012” is thus no longer significant in any specification, suggesting that the effect of the anticorruption campaign may not be as clear-cut.
Finally, the misinterpreted logarithm of area sold is also carried into CK’s Figure VI. Originally, CK find a consistent positive difference in the quantity of land purchased between princeling and nonprinceling firms, with the 95% confidence interval excluding 0 from 2007-2012, as is apparent in Figure 1(a). The quantity difference then “shrunk noticeably” after 2012, with the effect being statistically insignificant in 2013 and only marginally significant in 2015 and 2016 (p. 219). Reinterpreting the y-axis yields a smaller range of differences between princeling and nonprinceling land purchases, with the differences ranging from slightly below 0 to nearly 8,000 square meters when the outer bounds of the confidence interval are included.
In Figure VI without duplicates (illustrated in Figure 1(b)), the effect of the anticorruption campaign is less clear: the quantity difference in land purchased between princeling and nonprinceling firms drops in 2011 prior to the start of the anticorruption campaign, although the confidence intervals remain large. Further, the 95% confidence interval of the quantity difference between princelings and nonprincelings does not drop fully towards zero as it did in Figure 1(a), with the quantity difference remaining statistically significant after 2007. The effect is ultimately much less clean, again limiting arguments about the anticorruption campaign’s causal impact.
4 Discussion
The presence of duplicates and the misinterpreted logarithm of area raise important implications about CK’s original results, chiefly the size of the princeling effect in terms of square meters. While the princeling effect often remained statistically significant but shrunk when duplicates were removed (as apparent in Table 5), the economic significance is not fully clear: given the significant variation in firm parcel purchase volumes and sizes, a benefit of around 78,000 square meters (Table 5, column (1)) is extremely significant for a firm whose yearly purchase total per province is around the median of 26,114 square meters, but less so for one whose yearly purchase volume is over 4 million square meters (as is the case of the princeling with the largest land purchase total in a single year).
Importantly, these results also raise questions about underlying data quality; with the panel data used in CK’s Tables VI and XI having approximately 91% and 95% zeros, respectively, in the area column, effects are estimated from a small number of nonzero coefficients and are sensitive to duplicates. Given these consistency issues, taking the logarithm of the area would certainly afford more meaningful insight. Yet, when the actual logarithm of area is used (as described in the Appendix), the results are unstable, changing sign and direction across specifications.
Further, the causal implications of the anticorruption campaign are limited when duplicates are removed, with Figure 1(b) highlighting that the difference in princeling and nonprinceling land purchase volumes dropped before 2012. Similarly, Table 5 confirms that the coefficient for a transaction occurring after 2012 changes sign and is no longer statistically significant—although the variables representing central inspections and Xi-appointed officials remain significant—perhaps suggesting that the anticorruption campaign depends more on regional enforcement and signaling. Given these results as well as the instability of the analysis when the true logarithm is used, further investigation should be done to determine whether the anticorruption campaign truly plays a causal role in reducing princeling discounts.
5 Appendix A: Duplicate examples
Example of correct duplicates (two parcels that are right next to each other, having the same buyer, transaction month/year, and parcel area): princeling transaction IDs 3060341 and 3060345, which have electronic lookup numbers 6543222013B00600 and 6543222013B00595. The Transfer Announcements confirm that they are different parcels.
Incorrect duplicates:
1.
Duplicate listings with multiple parcels are assigned the same electronic lookup number, with the Transfer Announcement confirming there is only one parcel of the given size for sale.
Example: Princeling transaction IDs: 125957, 1527560.
Electronic lookup number (for both listings in the database): 2106242010B00016. -
2.
Electronic lookup numbers for multiple parcels are the same, although Transfer Announcements are not available to confirm whether the parcels are indeed the same.
Example: Princeling transaction IDs: 135826, 2991138.
Electronic lookup number (for both listings in the database): 2112232010B00057. -
3.
Duplicate listings have multiple electronic lookup numbers, but the Transfer Announcement suggests they refer to the same underlying parcel. In this case, one listing appears to be more complete than the other.
Example: Princeling transaction IDs: 3002177, 105512, 943888.
Electronic lookup numbers: 1525302010B00059, 1525302010B00076.
Note that in this case, there are three princeling transactions with the parcel features but only two listings in the Ministry of Land’s database, suggesting there are two erroneous duplicates. -
4.
Duplicate listings have high numbers of duplicates that are nonsensical and are sometimes wholly omitted from the database.
Example 1: 73 entries of a parcel with an area of 100 square meters purchased by the same buyer at the same price in the same month/year/province (all princelings).
Princeling transaction IDs from these listings include: 508039, 508046, 508048, 508049, 508063, ….
Example 2: 71 duplicates with an area of 289 square meters.
Princeling land transaction IDs: 321688, 321689, 321692, 321693, 321694, 321697, …. -
5.
Partially correct duplicates exist, wherein there are two parcels right next to each other, having the same buyer, transaction month/year, and area of the parcels. The Transfer Announcements suggest they are different parcels, but there are extra copies in the princeling database.
Example: Princeling transaction IDs: 2248276, 779393, and 1368322.
Electronic lookup numbers: 5301262010B00026, 5301262010B00019.
6 Appendix B: Misspecifying area/1,000,000 as log(area)
In CK’s replication data, the dataset price.dta contains the parcel-level transaction records described in section 2. By matching transaction records from the Land Transaction Monitoring System with CK’s dataset, I am able to discern the original areas of each parcel.
In their subsequent tables, CK aggregate this parcel-level data, summing the given purchases by each firm in each year (Table VI), by each firm in each year in each province (Table XI), and then by each year and province/municipality (Tables VIII, IX, and XII). At this step, the misinterpretation of area becomes apparent in a simple example about Firm 1.
In the year 2015, Firm 1 has only one transaction with a raw area of 17,795.3 square meters. Thus, when CK aggregate the total amount of area sold to each firm in each year, Firm 1’s total for 2015 is still just 17,795.3 square meters. Yet, in their dataset, the value that CK have for the reported is 0.01764, which is approximately 17,795.3/1,000,000. The slight variation in numbers seems to come from their calculation: they likely take the logarithm of the original area divided by 100. They then truncate/round it, exponentiate it, and divide it by 10,000.
| Firm ID | Year | Area | Area/1,000,000 | CK’s reported lnarea | Actual lnarea |
| 1 | 2015 | 17,795.3 | 0.017795 | 0.01764 | 9.78669 |
A second example with aggregation confirms that this is the case. For instance, firm 83 has 2 land transactions that occur in 2013. I include both of these transaction records, as well as the aggregated entry, in the two panels of Table 7.
| Panel A: Transaction-level data | |||||
| Firm ID | Year | CK transaction identifier | Electronic lookup no. | Area | CK’s reported lnarea |
| 83 | 2013 | 2086875 | 1310822013B00732 | 17481.17 | 5.17 |
| 83 | 2013 | 2090254 | 1310822013B00680 | 639.87 | 1.86 |
| Panel B: Data aggregated by firm | |||||
| Firm ID | Year | Area (Sum) | Area/1,000,000 | CK’s reported lnarea | Actual lnarea |
| 83 | 2013 | 18121.04 | 0.01812 | 0.01796 | 9.392 |
In these cases and all others, CK’s reported is approximately area/1,000,000, the slight difference again resulting from CK’s rounding during intermediate steps. In Section 3, I reinterpret CK’s main tables in light of the fact that area is scaled rather than logged; the tables recalculated with correct log values are included in Appendix C.
7 Appendix C: Calculations utilizing corrected log(area) values
In this Appendix, I conduct Chen and Kung’s calculations with the correct logarithm of area purchased and find, in many cases, substantially different results, with and without duplicates. In particular, the princeling effect changes its sign and direction, differing in cases with and without duplicates; in multiple instances, the results are qualitatively different when the correct logarithm of area value is used, even when calculated on the original data. Further, there is evidence of a sharp drop in the princeling advantage prior to the anticorruption campaign. As above, I regenerate Tables VI and XI as well as Figure VI, which are impacted by these errors. For this analysis, CK seek to analyze the differences in quantity purchased between princelings and nonprincelings overall, not just those firms who buy land in a given year—and presumably, princelings may be more likely to purchase land in a given year in the first place. Thus, I calculate log(area), not using a scaling factor, and following CK, I add 1 to the area value to “deal with observations of 0” prior to taking the log (p. 212).
For Table VI, CK originally interpret the regression coefficient as suggesting that princeling-connected firms purchased 0.2% (column (1), exp(0.002)) more land each year than their nonprinceling counterparts; they find the quantity of land purchased by firms connected to Politburo Standing Committee members (PSCM) is 3% higher (column (2), exp(0.001+0.031)) than that purchased by nonprincelings and that retirement has a negligible effect on the quantity of land purchased by princeling firms (column (3)). Yet, as highlighted in Panels A and B of Table 8, these results differ significantly from calculations that involve the correct logarithm of area values. In the baseline specification, princeling-connected firms purchase 12.4% more land each year than their nonprinceling counterparts. The other specifications are more complicated: column (2) suggests that princeling firms purchase 26.2% less land than their nonprinceling counterparts but that princeling firms with PSCM connections purchase a massive 179,084% more land (exp(0.304+7.795)1) than nonprinceling firms. Similarly, column (3) indicates that princelings purchase 6.1% less land than their nonprinceling counterparts but that princeling firms connected to retired officials purchase 344,511% more land (exp(0.063+8.208)1) than their nonprinceling
| Log of area of land purchased | ||||||
| Panel A: Original data | Panel B: Data without duplicates | |||||
| (1) | (2) | (3) | (1) | (2) | (3) | |
| Princeling firms | 0.117*** | -0.304*** | -0.063*** | 0.102*** | -0.287*** | -0.069*** |
| (0.013) | (0.011) | (0.012) | (0.013) | (0.011) | (0.011) | |
| Princeling firms*PSCM | 7.795*** | 7.208*** | ||||
| (0.082) | (0.082) | |||||
| Princeling firms*Retired | 8.208*** | 7.825*** | ||||
| (0.119) | (0.115) | |||||
| Adjusted | 0.044 | 0.046 | 0.045 | 0.044 | 0.046 | 0.045 |
peers. Such percentages are so large as to be nonsensical: with the median area of land purchased annually by a firm being 26,668 square meters, a 344,511% increase above this would be 91,900,739 square meters while a 179,084% increase would be 47,784,870 square meters—and these are well above the maximum annual firm purchase volume of around 14.7 million square meters.171717As above, this median statistic includes only the rows in the panel that have nonzero area values (pre-log); otherwise, the median would be 0. The mean is 6,250 square meters with zeros and 68,193 square meters excluding zeros, as a small number of firms have very high transaction amounts and skew the mean upwards. I thus find the median more accurately reflects the average firm’s yearly transaction volume but note that the princeling advantage would yield even more significant increases in area if the mean were used instead. These specifications are also interesting because columns (2) and (3) suggest that regular princelings purchase less land than nonprincelings, but princelings with PSCM and retired official connections purchase significantly more, reversing CK’s original conclusions—and the regressions calculated without duplicates (Panel B, columns (1)-(3)) further echo these results.
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign. Table XI thus regresses the logarithm of land area purchased on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurred in a province wherein Xi replaced the party secretary (column (3)). CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1%; after 2012, princeling land purchases dropped by more than 2.3% (and by 29% in their purchase quantity premium compared to nonprincelings). With the corrected log area, as shown in Panel A of Table 9, the regressions suggest princeling firms purchased 244,694% (exp(7.803)1) more land than their nonprinceling counterparts prior to 2012, and after 2012, their land purchases increase by 52,112% (exp(7.803)exp(7.803+0.193)), or a 21.3% increase (() in their purchase quantity premium compared to nonprincelings.
| Log of area of land purchased | ||||||||
| Panel A: Original data | Panel B: Data without duplicates | |||||||
| (1) | (2) | (3) | (4) | (1) | (2) | (3) | (4) | |
| Princeling firms | 7.803*** | 7.938*** | 7.905*** | 7.800*** | 7.226*** | 7.471*** | 7.401*** | 7.223*** |
| (0.205) | (0.189) | (0.205) | (0.207) | (0.211) | (0.198) | (0.216) | (0.213) | |
| Princeling firms | 0.193 | 0.770*** | 0.660*** | 1.193*** | ||||
| Transactions after 2012 | (0.150) | (0.154) | (0.148) | (0.156) | ||||
| Central inspection | 9.014*** | 5.523*** | 8.935*** | 5.461*** | ||||
| (0.071) | (0.895) | (0.085) | (0.889) | |||||
| Princeling firms | -8.468*** | -5.495*** | -8.033*** | -5.428*** | ||||
| Central inspection | (0.199) | (0.925) | (0.201) | (0.916) | ||||
| Xi-appointed officials | 9.079*** | 7.307*** | 9.024*** | 7.272*** | ||||
| (0.115) | (0.399) | (0.114) | (0.394) | |||||
| Princeling firms | -8.260*** | -7.040*** | -7.785*** | -6.939*** | ||||
| Xi-appointed officials | (0.205) | (0.439) | (0.214) | (0.433) | ||||
| Adjusted | 0.094 | 0.198 | 0.250 | 0.283 | 0.094 | 0.205 | 0.262 | 0.297 |
For the interacted central inspection equation (column (2)), the estimation suggests that the land sold to princeling firms increased by 203,501% (exp(7.938)exp(7.938+9.0148.468)), or that their purchase quantity premium increased by 72.7% compared to nonprinceling firms (). These percentages are so large as to be wholly nonsensical: with the median total purchased by a firm in a given year and province being 26,114 square meters, these numbers suggest the average princeling advantage would be 63,925,369 square meters (per column (1)’s 244,694% growth)—a value far above the maximum annual land purchase by a firm.181818As noted in the main text, many firms purchase land infrequently, and 95.4% of the observations have zero square meters purchased in a given province/year. The reported median includes only those nonzero transactions (as the median would otherwise be 0). The mean is 3,076 square meters when including rows with zero square meters sold and 66,423 square meters when excluding zeros. Likewise, in CK’s original analysis, they find that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant in column (4) when the dummy variables for central inspection and Xi-appointed officials are included, but when recalculated, this is no longer the case.
When duplicate rows are removed, as highlighted in Panel B of Table 9, the coefficients are still quite large: Panel B’s column (1), for instance, indicates that princeling firms purchased more land than their nonprinceling counterparts by 137,371% (exp(7.226)1). After 2012, their land purchases increase by 128,507% (exp(7.226)exp(7.226+0.660)), which is a 93.5% increase () in their purchase quantity premium relative to nonprincelings. Yet, due to the extremely large percentages of princeling advantage in purchase quantity premium, the results are difficult to interpret and nonsensical, likely reflecting high variability in parcel sizes, the need for further controls, the relatively small number of princeling firms in the overall data, and the sensitivity resulting from adding 1 to area prior to logging it. These results are also interesting because they run counter to that from the recalculated Table VI, which suggested princeling firms—in the presence of interacted terms—purchased less land than their nonprinceling counterparts, confirming these data and estimation issues.
In their original results for Figure VI, CK find that the consistent positive difference in the quantity of land purchased between princeling and nonprinceling firms “shrunk noticeably” after 2012, with the effect being statistically insignificant in 2013 and only being “marginally significant in 2016” (p. 219). Yet, when generating Figure VI with the corrected logarithm of area (Figure 2(a)), the quantity difference between princeling and nonprinceling land purchases drops toward zero in 2011, prior to the anticorruption campaign, no longer positive or statistically significant. Likewise, the quantity difference and its 95% confidence interval have dropped below 0 in 2012, suggesting princelings purchase less land than their nonprinceling counterparts, and while the drop steepens in 2013 before eventually ticking back to zero in 2016, it seems that the declining trend began well before the anticorruption campaign—the effect is thus much less clean, limiting arguments about the causal impact of the anticorruption campaign.
This behavior is further apparent when the case without duplicates is examined (Figure 2(b)). While princelings purchase more land than nonprincelings from 2008-2011, the 95% confidence intervals drop towards zero in 2011 and include zero for all subsequent years, except 2014. This behavior again suggests that the quantity difference between princeling and nonprinceling land purchases dropped prior to the anticorruption campaign and remained at or slightly near zero for the bulk of its duration.